OpenStudy (anonymous):

5u+2v=-16, 3u+v=-7 solve by elimination

OpenStudy (anonymous):

which variable do you want to get rid of first: 5u+2v=-16 3u+v=-7

OpenStudy (anonymous):

u

OpenStudy (anonymous):

why the v seems easier

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

how

OpenStudy (anonymous):

so to get rid of the v well multiply that equation by -2 5u+2v=-16 (-2)[3u+v]=-7(-2) +__________________

OpenStudy (anonymous):

2*3

OpenStudy (anonymous):

ya so your problem will look like this now 5u+2v=-16 -6u-2v=14 +__________________

OpenStudy (anonymous):

ok what is next

OpenStudy (anonymous):

add them up

OpenStudy (anonymous):

11u,4v

OpenStudy (anonymous):

5u+(-6u)=-u

OpenStudy (anonymous):

-2v+2v=what

OpenStudy (anonymous):

4v

OpenStudy (anonymous):

-2v+2v=0

OpenStudy (anonymous):

now whats -16+14

OpenStudy (anonymous):

30

OpenStudy (anonymous):

why are you not including the negative in your adding

OpenStudy (anonymous):

sorry 28

OpenStudy (anonymous):

whats 14-16

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

correct so what -16+14

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

how are you getting that

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

adding -16+14

OpenStudy (anonymous):

you said 14-16=-2 so isnt -16+14 the same thing

OpenStudy (anonymous):

am I subtracting or adding

OpenStudy (anonymous):

what i wrote is the same thing i just switched places, its the same problem

OpenStudy (anonymous):

ok got it is = to 2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

now what do we do

OpenStudy (anonymous):

alryte so we have: -u=-2 we want the u to be positive so well divide both sides by negative 1 since -u is the same thing as (-1)u so -u/-1=-2/-1

OpenStudy (anonymous):

ok whats next

OpenStudy (anonymous):

what did u get

OpenStudy (anonymous):

-u/-1=-2/-1

OpenStudy (anonymous):

ya thats the problem what was the result tho